1.Three point charges of +16uC are placed at corners of an equilateral triangle of sides of length 2 cm. Calculate potential energy of the system measured with respect to zero when the charges are at infinite distance apart. 2.Two point charges of +20uC are placed 2cm apart. Calculate electric potential and electric field at mid-point of the line joining two charges
1
Expert's answer
2021-10-28T08:51:54-0400

1)

"\varphi=3\frac {kq}r=21.6~MV,"

2)

"\varphi=2\frac {kq}{\frac r2}=\frac{4kq}r=36~MV,"

"E=E_1-E_2=\frac{kq}{(\frac r2)^2}-\frac{kq}{(\frac r2)^2}=0."